# A proof of Alexander’s conjecture on an inequality of Cassels

### v1

@@ 1. Introduction
Let $z_{1},\dots,z_{n}$ be complex numbers in the unit disc. A classical inequality, going back to Schur [1], states that
@@
with equality if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
@@
Let $\rho>1$ and let $z_{1},\dots,z_{n}$ be complex numbers in the disc $\{z:|z|\leq\rho\}$. Suppose
@@
with equality if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
@@
In a recent work, Dubickas [4] revisited Cassels’ inequality by observing that, for $z_{j}=\rho\,\omega_{j}$ with $|\omega_{j}|=1$, one has
@@
He conjectured that the elementary symmetric functions of the quantities $|\omega_{j}-\omega_{k}|^{2}$ attain their maximum for the vertices of a regular $n$-gon inscribed in the unit circle, and proved this for the elementary symmetric functions of degree 11 through 44. His conjecture implies Alexander’s.
@@ 2. Main result
Let $\rho>1$ and let $z_{1},\dots,z_{n}$ be complex numbers in the disc $\{z:|z|\leq\rho\}$. Then the following inequality holds:
@@
Equality is attained if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the circle of radius $\rho$.
@@ 3. An auxiliary inequality
Let $z_{1},\dots,z_{n}$ on the unit disc and denote $\Lambda=(-1)^{n}\prod_{j=1}^{n}z_{j}$. Then we have the following inequality:
@@
Equality is attained if and only if the $z_{j}$ are the $n$ roots of the polynomial $\zeta^{n}+\Lambda$.
@@
Let $\omega_{1},\dots,\omega_{n}$ on the unit circle and $a\in(0,1)$. Then the following inequality holds:
@@
Equality is attained if and only if the $\omega_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
4. Proof of the main theorem
Proof of Theorem 2.
@@ 4. Proof of the main theorem
Setting $z_{j}=\rho\,\omega_{j}$ with $|\omega_{j}|=1$, (4) is equivalent to
@@
We fix $\omega_{1},\dots,\omega_{n}$ in the unit circle and we define, for $a\in[0,1)$,
$g_{\omega}(a)=n\ln(1-a^{n})-\sum_{j,k}\ln|a-\overline{\omega_{j}}\omega_{k}|.$
Using $\frac{d}{da}\log|a-\zeta|=\Re\frac{1}{a-\zeta}$ and the fact that $\sum_{j,k}\!\left(\frac{1}{a-\overline{\omega_{j}}\omega_{k}}\right)$ is real, we compute
@@
Now assuming that $z_{1},\dots,z_{n}$ realise equality in (4), they also realise equality in (7). Hence $z_{j}=\rho\omega_{j}$ with $|\omega_{j}|=1$ for $j=1,\dots,n$ and $g_{\omega}(\rho^{-2})=0$. Since $g_{\omega}$ is nondecreasing and $g_{\omega}(0)=g_{\omega}(\rho^{-2})=0$, it follows that $g_{\omega}$ is constant on $0,\rho^{-2}]$, hence $g_{\omega}^{\prime}(\rho^{-2})=0$, that is,
@@
By Corollary 4, the $\omega_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
Conversely, if the $z_{j}$ are the vertices of a regular $n$-gon, then
$\prod_{j=1}^{n}(\zeta-z_{j})=\zeta^{n}+\Lambda,\qquad|\Lambda|=\rho^{n}.$
@@ 5. Proof of Theorem 3
Proof.
Fix $z_{1},\dots,z_{n}$ on the unit disc and define
$f(\zeta):=\sum_{j=1}^{n}\frac{1}{1-\overline{z_{j}}\zeta},$
and the finite Blaschke product
$B(\zeta):=\prod_{j=1}^{n}\frac{\zeta-z_{j}}{1-\overline{z_{j}}\zeta}.$
Both are holomorphic in a neighborhood of the unit disc.
Expand
$|f(e^{it})|^{2}=\sum_{j,k}\frac{1}{(1-\overline{z_{j}}e^{it})(1-z_{k}e^{-it})}.$
Then
$\frac{1}{2\pi}\int_{0}^{2\pi}\frac{dt}{(1-\overline{z_{j}}e^{it})(1-z_{k}e^{-it})}=\frac{1}{2\pi i}\oint_{|\zeta|=1}\frac{d\zeta}{(1-\overline{z_{j}}\zeta)(\zeta-z_{k})}=\frac{1}{1-\overline{z_{j}}z_{k}},$
by the Cauchy integral formula. Summing over $j,k$ gives
(9) $\sum_{j,k}\frac{1}{1-\overline{z_{j}}z_{k}}=\frac{1}{2\pi}\int_{0}^{2\pi}|f(e^{it})|^{2}dt.$
For all $t\in[0,2\pi]$, since $|B(e^{it})|=1$, we have
@@
Integrating and using Cauchy integral formula $\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,dt=B(0)=\Lambda$ yields
@@
Using twice again Cauchy integral formula,
(11) $n=f(0)=\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\,dt,$
and
(12) $\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,\overline{f}(e^{it})\,dt=\sum_{j=1}^{n}\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,\frac{e^{it}}{e^{it}-z_{j}}\,dt=\sum_{j=1}^{n}B(z_{j})=0.$
Now, we multiply the complex conjugate of (12) by $\Lambda$ and we subtract it from (11) to find
(13) $\begin{split}n&=\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\,dt=\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\left(1-\Lambda\overline{B}(e^{it})\right)\,dt.\end{split}$
Applying Cauchy–Schwarz inequality to the integral in (13) and using (9) and (10) gives the desired inequality
(14) $\begin{split}n^{2}&\leq\Big(\frac{1}{2\pi}\int_{0}^{2\pi}|1-\Lambda\overline{B(e^{it})}|^{2}\,dt\Big)\Big(\frac{1}{2\pi}\int_{0}^{2\pi}|f(e^{it})|^{2}\,dt\Big)\\ &=(1-|\Lambda|^{2})\sum_{j,k}\frac{1}{1-\overline{z_{j}}z_{k}}.\end{split}$
Moreover, equality in (14) holds if and only if
@@
$\begin{split}P(\zeta)&=\prod_{j=1}^{n}(\zeta-z_{j})=\sum_{m=0}^{n}(-1)^{\,n-m}e_{n-m}\,\zeta^{m}\\ Q(\zeta)&=\prod_{j=1}^{n}(1-\overline{z_{j}}\zeta)=\sum_{m=0}^{n}(-1)^{m}\overline{e_{m}}\,\zeta^{m},\\ R(\zeta)&=\sum_{j=1}^{n}\prod_{k\neq j}(1-\overline{z_{k}}\zeta)=\sum_{m=0}^{n-1}(-1)^{m}(n-m)\,\overline{e_{m}}\,\zeta^{m},\end{split}$
@@
Identifying the coefficients in front of $\zeta^{m}$ for $m=1,\dots,n-1$, we get
@@
$P(\zeta)=\zeta^{n}+(-1)^{n}e_{n}=\zeta^{n}+\Lambda.$
Therefore $z_{1},\dots,z_{n}$ are precisely the roots of $\zeta^{n}+\Lambda$.
Conversely, if $z_{1},\dots,z_{n}$ are the roots of $\zeta^{n}+\Lambda$, then $e_{m}=0$ for $1\leq m\leq n-1$ and the above coefficient comparison can be reversed to obtain (15), hence equality in (14). ∎

### v2

@@ 1. Introduction
We will use the following notations for the unit disc and the unit circle :
$\mathbb{D}=\{z\in\mathbb{C}:|z|<1\},\ \ \overline{\mathbb{D}}=\{z\in\mathbb{C}:|z|\leq 1\},\ \ \mathbb{T}=\{z\in\mathbb{C}:|z|=1\}.$
Let $z_{1},\dots,z_{n}$ be in $\mathbb{D}$. A classical inequality, going back to Schur [1], states that
@@
with equality if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in $\mathbb{T}$.
@@
Let $\rho>1$ and let $z_{1},\dots,z_{n}\in\{z\in\mathbb{C}:|z|\leq\rho\}$. Suppose
@@
with equality if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the circle $\{z\in\mathbb{C}:|z|=\rho\}$.
@@
In a recent work, Dubickas [4] revisited Cassels’ inequality by observing that, for $z_{j}=\rho\,\omega_{j}$ with $\omega_{j}\in\mathbb{T}$ ($j=1,\dots,n$), one has
@@
He conjectured that the elementary symmetric functions of the quantities $\{|\omega_{j}-\omega_{k}|^{2},\ j<k\}$ attain their maximum for the vertices of a regular $n$-gon inscribed in $\mathbb{T}$, and proved this for the elementary symmetric functions of degree 11 through 44. His conjecture implies Alexander’s.
@@ 2. Main result
Let $\rho>1$ and let $z_{1},\dots,z_{n}\in\{z\in\mathbb{C}:|z|\leq\rho\}$. Then the following inequality holds:
@@
Equality is attained if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the circle $\{z\in\mathbb{C}:|z|=\rho\}$.
@@ 3. An auxiliary inequality
Let $z_{1},\dots,z_{n}\in\mathbb{D}$ and denote $\Lambda=(-1)^{n}\prod_{j=1}^{n}z_{j}$. Then we have the following inequality:
@@
Equality is attained if and only if the $z_{j}$ are the $n$ roots of the polynomial $P(z)=z^{n}+\Lambda$.
@@
Let $\omega_{1},\dots,\omega_{n}\in\mathbb{T}$ and $a\in(0,1)$ . Then the following inequality holds:
@@
Equality is attained if and only if the $\omega_{j}$ are the vertices of a regular $n$-gon inscribed in $\mathbb{T}$.
Proof.
Apply Theorem 3 to $z_{j}=\sqrt{a}\,\omega_{j}\in\mathbb{D}$ $(j=1,\dots,n)$. Then
$1-\overline{z_{j}}z_{k}=1-a\,\overline{\omega_{j}}\omega_{k},\qquad|Z|^{2}=\Bigl|\prod_{j=1}^{n}z_{j}\Bigr|^{2}=a^{n}.$
This gives (6). The equality case follows from that of Theorem 3. ∎
4. Proof Theorem 2
@@
Setting $z_{j}=\rho\,\omega_{j}$ with $\omega_{j}\in\mathbb{T}$ ($j=1,\dots,n)$, (4) is equivalent to
@@
We fix $\omega_{1},\dots,\omega_{n}$ on $\mathbb{T}$ and we define, for $a\in[0,1)$,
$g_{\omega}(a)=n\log(1-a^{n})-\sum_{j,k}\log|a-\overline{\omega_{j}}\omega_{k}|.$
Using $\frac{d}{da}\log|a-z|=\Re\frac{1}{a-z}$ and the fact that $\sum_{j,k}\!\left(\frac{1}{a-\overline{\omega_{j}}\omega_{k}}\right)$ is real, we compute
@@
Now assuming that $z_{1},\dots,z_{n}$ realize equality in (4), they also realize equality in (7). Hence $z_{j}=\rho\omega_{j}$ with $\omega_{j}\in\mathbb{T}$ ($j=1,\dots,n$) and $g_{\omega}(\rho^{-2})=0$. Since $g_{\omega}$ is nondecreasing and $g_{\omega}(0)=g_{\omega}(\rho^{-2})=0$, it follows that $g_{\omega}$ is constant on $0,\rho^{-2}]$, hence $g_{\omega}^{\prime}(\rho^{-2})=0$, that is,
@@
By Corollary 4, the $\omega_{j}$ are the vertices of a regular $n$-gon inscribed in $\mathbb{T}$.
Conversely, if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in $\{z\in\mathbb{C}:|z|=\rho\}$, then
$P(z)=\prod_{j=1}^{n}(z-z_{j})=z^{n}+\Lambda,\qquad|\Lambda|=\rho^{n}.$
@@ 5. Proof of Theorem 3
Recall that the Cauchy integral formula gives, for a holomorphic function $F$ in a neighborhood of $\overline{\mathbb{D}}$ and for $\zeta\in\mathbb{D}$,
(9) $\frac{1}{2\pi}\int_{0}^{2\pi}\frac{e^{it}F(e^{it})}{e^{it}-\zeta}\,dt=\frac{1}{2\pi i}\int_{|z|=1}\frac{F(z)}{z-\zeta}\,dz=F(\zeta).$
Fix $z_{1},\dots,z_{n}$ in $\mathbb{D}$ and define the folowing functions, holomorphic in a neighborhood of $\overline{\mathbb{D}}$:
$f(z)=\sum_{j=1}^{n}\frac{1}{1-\overline{z_{j}}z},\ \ B(z)=\prod_{j=1}^{n}\frac{z-z_{j}}{1-\overline{z_{j}}z}.$
Note that for all $t\in[0,2\pi]$,
$|f(e^{it})|^{2}=\sum_{j,k}\frac{e^{it}}{(1-\overline{z_{j}}e^{it})(e^{it}-z_{k})}.$
By (9), we have
(10) $\frac{1}{2\pi}\int_{0}^{2\pi}|f(e^{it})|^{2}\,dt=\sum_{j,k}\frac{1}{2\pi}\int_{0}^{2\pi}\frac{e^{it}}{(1-\overline{z_{j}}e^{it})(e^{it}-z_{k})}\,dt=\sum_{j,k}\frac{1}{1-\overline{z_{j}}z_{k}},$
(11) $\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,\overline{f}(e^{it})\,dt=\sum_{j=1}^{n}\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,\frac{e^{it}}{e^{it}-z_{j}}\,dt=\sum_{j=1}^{n}B(z_{j})=0,$
(12) $\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,dt=B(0)=\Lambda,$
and
(13) $\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\,dt=f(0)=n.$
We multiply the complex conjugate of (11) by $\Lambda$ and we subtract it from (13) to find
(14) $\begin{split}n&=\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\left(1-\Lambda\overline{B}(e^{it})\right)\,dt.\end{split}$
Besides, for all $t\in[0,2\pi]$, since $|B(e^{it})|=1$, we have
@@
Integrating and using (12) yields
@@
Now (10), (14), (15) and Cauchy–Schwarz applied to the integral in (14) gives the desired inequality:
(16) $\begin{split}n^{2}&\leq(1-|\Lambda|^{2})\sum_{j,k}\frac{1}{1-\overline{z_{j}}z_{k}}.\end{split}$
We now turn to the equality case. Equality in (16) holds if and only if
@@
$\begin{split}P(z)&=\prod_{j=1}^{n}(z-z_{j})=\sum_{m=0}^{n}(-1)^{\,n-m}e_{n-m}\,z^{m}\\ Q(z)&=\prod_{j=1}^{n}(1-\overline{z_{j}}z)=\sum_{m=0}^{n}(-1)^{m}\overline{e_{m}}\,z^{m},\\ R(z)&=\sum_{j=1}^{n}\prod_{k\neq j}(1-\overline{z_{k}}z)=\sum_{m=0}^{n-1}(-1)^{m}(n-m)\,\overline{e_{m}}\,z^{m},\end{split}$
@@
Identifying the coefficients in front of $z^{m}$ for $m=1,\dots,n-1$, we get
@@
$P(z)=z^{n}+(-1)^{n}e_{n}=z^{n}+\Lambda.$
Therefore $z_{1},\dots,z_{n}$ are precisely the roots of $z^{n}+\Lambda$.
Conversely, if $z_{1},\dots,z_{n}$ are the roots of $z^{n}+\Lambda$, then $e_{m}=0$ for $1\leq m\leq n-1$ and the above coefficient comparison can be reversed to obtain (17), hence equality in (16).